claire is one year older than her sister nancy. nancy is three years older than her sister sheila. sheila is one year older than her brother david. david is five years older than his brother mike. the sum of the five siblings' ages is 120 years. find the age of each.
Answer by ewatrrr(24785)

 
Hi,
Let x represent Mike's age...then
David is(x+5)
sheila is (x+6)
nancy is (x+9)
claire is (x+10)
Question states***
x + (x+5)+(x+6)+(x+9)+(x+10) = 120
5x +30 = 120
5x = 90
x = 18, Mike's age. David 23, Sheila 24, Nancy 27 and Claire 28
CHECKING our Answer***
18 + 23 + 24 + 27 + 28 = 120